What two numbers when added equal negative 63, and when the second number is subtracted from the first they equal negative 41.
Answer by Maths68(1474)


Let
First Number = x
Second Number = y
x+y=-63........(1)
x-y=41.........(2)
Add (1) to (2)
x+y=-63
x-y=41
-------------
2x=-22
2x/2=-22/2
x=-11
Put the value of x in (1)
x+y=-63
-11+y=-63
y=-63+11
y=-52
First Number = x = -11
Second Number = y = -52


Check
==========
Two numbers when added equal negative 63
-11-52=-63
-63=-63
The second number is subtracted from the first they equal negative 41.
-11-(-52)=41
-11+52=41
41=41